A bourbon that is 54-proof is 27% alcohol by volume while one that is 81-proof is 40.5% alcohol. How many liters of 54-proof bourbon must be mixed with 1.0-L of 81-proof bourbon to produce a 66-proof bourbon? Give your answer to three decimal places.

Answer by stanbon(75887)

A bourbon that is 54-proof is 27% alcohol by volume while one that is 81-proof is 40.5% alcohol. How many liters of 54-proof bourbon must be mixed with 1.0-L of 81-proof bourbon to produce a 66-proof bourbon? Give your answer to three decimal places.
------
Equation:
alcohol + alcohol = alcohol
0.27x + 0.405*1 = 0.33(x+1)
--------
270x + 405 = 330x + 33
------
60x = 372
x = 6.200 L (amt of 54 proof needed)
================
Cheers,
Stan H.